OpenStudy (anonymous):

find the standard equation, center, and radius of 6x^2+6y^2+24x+12y=-6 Show work.

OpenStudy (anonymous):

you know how to complete the square?

OpenStudy (anonymous):

kinda, i dont understand how to do it, so you could say no

OpenStudy (anonymous):

first group the terms: 6x^2+24x+6y^2+12y=-6

OpenStudy (anonymous):

now complete the square for each variable

OpenStudy (anonymous):

how?

myininaya (myininaya):

\[6(x^2+4x)+6(y^2+2y)=-6\] divde both sides by 6 \[x^2+4x+y^2+2y=-1\] \[(x^2+4x+(\frac{4}{2})^2)+(y^2+2y+(\frac{2}{2})^2)=-1+(\frac{4}{2})^2+(\frac{2}{2})^2\]

myininaya (myininaya):

\[(x^2+4x+2^2)+(y^2+2y+1^2)=-1+4+1\] \[(x+2)^2+(y+1)^2=4\]

OpenStudy (anonymous):

 you can complete the rest i hope or someone else will :)